I have been given a function table and have to come up with a formula. The input (x)is sometimes given, and the output(y) is sometimes given. This is what I have: (1, 8), (2, 14),(3,y), (x, 24), (x, 28), (6,y). I noticed the difference between the output goes from +6 to +4. I think the value of y when x=3 is 18. I also think the values of x increase by one each time. I have come up with two formulas: If x is even, the formula is (x * 5) + 4. If x is odd, the formula is (x * 5) + 3.Is it valid to have two formulas? Or do I have to show the odd and even numbers some other way?
Answer by stanbon(48558)

I have been given a function table and have to come up with a formula. The input (x)is sometimes given, and the output(y) is sometimes given. This is what I have: (1, 8), (2, 14),(3,y), (x, 24), (x, 28), (6,y).
----------------
Looking at the 1st two points you have a slope of 6.
If x was 0, y would be 2.
EQUATION: y=6x+2
For (3,y), y=6*3+2=20
For (x,24), 24=6x+2, 22=6x, x=11/3
For (x,28), 28=6x+2, 6x=26, x=13/3
For (6,y), y=6*6+2 = 38
=============
Cheers,
Stan H.